2concentric spherical cells of radius a and b (a>b) are having uniformly distributed charges q1 and q2 resp.find out total energy

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

Potential inside all points of the sphere of radius a due to charge over its surface=V=\dfrac{kq_1}{a}

All the charges residing on the surface of shell with radius b are under the same potential. Hence potential energy of the system=q_2\times V=\dfrac{kq_1q_2}{a}